# Mapping Cone Connections and their Yang-Mills Functional

### v1

@@ 1 Introduction
For the physical motivation, the cone Yang-Mills functional can be considered as the dimensional reduction of the standard Yang-Mills functional over the $S^{1}$ fiber of a circle bundle, $\pi:X\to M$. In the generic case where the circle bundle $X$ is non-trivial, the twisting of the $S^{1}$ fiber over $M$ can be interpreted as sourced by a Kaluza-Klein monopole with a non-zero two-form flux $\zeta$ (mathematically, the Euler class of the circle bundle). On $X$, there is a global angular one-form $\theta$ which locally, can be expressed by $\theta=dz+a$ where $z$ is the $S^{1}$ coordinate and $a$ is the local $U(1)$ gauge field. Then the two-form flux, $\zeta=d\theta=d(dz+a)=da$, is the curvature of the $U(1)$ gauge field which measures the twist of the $S^{1}$ fiber.
Starting with the circle bundle $\pi:X\to M$, we dimensionally reduce the Yang-Mills functional by taking the metric to be
@@
and further require that the local connection one-form (gauge field) $\mathcal{A}$ on $X$ be invariant under translation of the circle fiber. Such a connection form on $X$ can be expressed as
@@
Here, both $A$ and $B$ takes values on $Ad\,P$, and $B$ represents the component of the connection in the circle direction. Since $\mathcal{A}$ is $S^{1}$ invariant, both $A$ and $B$ only have dependence on the coordinates of $M$. The curvature two-form (gauge field strengths) of such a connection one-form on $X$ then has the following expression:
@@
where we have used $d\theta=\zeta$. Comparing with the cone Yang-Mills functional in (1), we see that $S_{cYM}=\|\mathcal{F}_{\mathcal{A}}\|^{2}\,$, and therefore, it is just the Yang-Mills functional on $X$ dimensionally reduced to $M$ over the fiber circle. And we can also interpret the cone Yang-Mills Euler-Lagrange equations (1.2)-(1.3) as the dimensionally-reduced Yang-Mills equations on $X$.
@@
where we have used a simplified notation, e.g. $F_{A}\wedge*F_{A}=|F_{A}|^{2}dV$. The action thus consists of a Yang-Mills part, an adjoint-valued scalar with space-dependent “mass” $|\zeta|$ (which in general is not a constant over $M$), and an interaction term between $A$ and $B$. Roughly then, the Euler-Lagrange equations of motion (1.2)-(1.3) can be thought of as a mixing of the Yang-Mills equations with a massive Klein-Gordon scalar plus an interaction term.
The cone Yang-Mills functional also has a mathematical motivation. In fact, the functional arises naturally from its relation with a mapping cone complex which we will provide an explanation here. (For a reference on the general mapping cone complex, see for example [Weibel, CTT].)
@@
Since the grading of the two components of $\operatorname{Cone}^{k}(\zeta)$ in (1.8) are different, it is useful to introduce a formal one-form $\theta$ to supplement the second component so that we can express a cone form as a sum with both components having the same total degree, i.e.
@@
Additionally, requiring the property that $d\theta=\zeta$, the expression (1.9) allows us to interpret the cone differential $d_{C}$ simply as the exterior derivative:
@@
where $\mathcal{A}=A+\theta\,B\,$, with $A\in\Omega^{1}(M,Ad\,P)$ and $B\in\Omega^{0}(M,Ad\,P)$. The cone curvature then takes the form
@@
The cone Yang-Mills functional of (1) is then just the normed square of the cone curvature, $\|\mathcal{F}_{\mathcal{A}}\|^{2}$. We note that from this mapping cone perspective, the closed two-form $\zeta\in\Omega^{2}(M)$ need not be an element of $H^{2}(M,\mathbb{Z})$, which would be required if we consider $\zeta$ to be Euler class of a circle bundle.
In this paper, we take an important first step in understanding cone Yang-Mills solutions. We will show that a subset of the solutions consist of Yang-Mills solutions. For instance, in dimension two when $M$ is compact and $\zeta$ is taken to be the volume form, the above cone-flat condition (1.11) implies exactly the two-dimensional Yang-Mills condition
@@
On the other hand, it should be evident that the space of cone Yang-Mills solutions $(A,B)$ is generally much richer than that of Yang-Mills solutions. As we will show in explicit examples, not all Yang-Mills connection $A$ can be paired with a scalar $B$ to form a cone Yang-Mills solutions. Conversely, there are also cone Yang-Mills solutions $(A,B)$ where $A$ is not Yang-Mills. Furthermore, given a cone Yang-Mills solutions $(A,B)$, there may be other scalars $B^{\prime}$ such that $(A,B^{\prime})$ remain cone Yang-Mills.
When $M$ is three-dimensional, we obtain a Bogomolny-type condition, that gives a sufficient condition on $(A,B)$ to be a cone Yang-Mills solution. We derive such a condition by dimensionally reducing the four-dimensional (anti-)self-dual Yang-Mills condition for $\mathcal{F}_{\mathcal{A}}$, expressed in terms of (1.10). The resulting three-dimensional condition is
$F_{A}+\zeta B=\pm*d_{A}B$ (1.13)
Notice that if we set $\zeta=0$, the above equation is just the Bogomolny monopole equation. Similar to the (anti-)self-dual Yang-Mills condition in four dimensions, (1.13) is a first-order condition whose solutions solve (1.2)-(1.3) in three dimensions. This duality condition will assist us in finding non-abelian, cone Yang-Mills solutions. Indeed, we shall give an explicit $SU(2)$ solution of (1.13) that comes from dimensionally reducing the four-dimensional Taub-NUT gravitational instanton solution.
@@
This paper is organized as follows. In Section 2, we give general properties of the cone Yang-Mills equations and their solutions. In Section 3, we describe special cone Yang-Mills solutions in dimension two and three. In Section 4, we study the cone-flat connections and prove the classification of cone-flat bundles of Theorem 1.1.
@@ 2.1 Preliminaries
Let $M$ be a smooth manifold and $P$ be a principal bundle over $M$. We consider the associated adjoint bundle $Ad\,P$ equipped with an inner product that is invaraint under the adjoint action. Hence, for arbitrary $x,y,z\in\Omega^{*}(M,Ad\,P)$,
@@
Having defined the inner product on $\operatorname{Cone}^{*}(\zeta)(M,Ad\,P)$, we write down the Yang-Mills functional associated with the cone connection. Given a connection $A\in\Omega^{1}(M,Ad\,P)$ and a section $B\in\Omega^{0}(M,Ad\,P)$, we define the cone covariant derivative to be $D_{\mathcal{C}}=d_{A}+\theta B$. Locally, if we write the usual covariant derivative as $d_{A}=d+A$, then $D_{\mathcal{C}}=d+A+\theta B$ and the cone curvature takes the form
$\mathcal{F}=D_{\mathcal{C}}D_{\mathcal{C}}$ $=d(A+\theta B)+(A+\theta B)\wedge(A+\theta B)$
@@
$S_{cY\!M}(A+\theta B)$ $=\|\mathcal{F}\|_{\mathcal{C}}^{2}=\|(F_{A}+\zeta B)-\theta\,d_{A}B\|^{2}_{C}$
@@
$=\|\mathcal{F}\|_{\mathcal{C}}^{2}+2t\big{(}\langle d_{A}^{*}(F_{A}+\zeta B)+[B,d_{A}B],\eta\rangle+\langle\zeta^{*}(F_{A}+\zeta B)+d_{A}^{*}d_{A}B,\xi\rangle\big{)}+o(t)$
@@ 2.2 Elliptic property of the cone Yang-Mills solutions
We can write (2.13)-(2.14) more simply in terms of $D_{\mathcal{C}}$, the cone covariant derivative. Locally, if $d_{A}=d+A$, then $D_{\mathcal{C}}=d+A+\theta B$ and
@@
The last equation follows from the relation $(-1)^{m+1}*[\xi,*d_{A}B]=[\xi,d_{A}B]$. This can be seen by writing $d_{A}B=\sum\mu_{i}\otimes\alpha_{i}$ and $\xi=\sum\xi_{i}\otimes\alpha_{i}$ where $\mu_{i}\in\Omega^{1}(M)$, $\xi_{i}\in\Omega^{0}(M)$ and $\{\alpha_{i}\}$ is a basis of $\mathfrak{g}$. Then
@@
2.3 Comparision with Yang-Mills solutions
@@
Hence, we see that cone Yang-Mills solutions are a subset of Yang-Mills solutions satisfying additionally the second condition of (2.20).
In particular, when $M$ is a Kähler manifold and $\zeta=\omega$ is taken to be the Kähler metric, a hermitian Yang-Mills connection which satisfies the conditions
@@
is also a critical point of the cone Yang-Mills functional satisfying (2.20) with $B=0$. This is because $\omega^{*}F_{A}=0$ is equivalent to $\omega^{n-1}\wedge F_{A}=0$ for a Kähler metric.
@@
When $\zeta=0$, $\zeta^{*}$ is a zero map. The second equation of cone Yang-Mills condition (2.10) becomes just $d_{A}^{*}d_{A}B=0$, which implies $d_{A}B=0$ on a compact manifold. The first equation (2.9) then simplifies to $d_{A}^{*}F_{A}=0$. Hence, $A$ must be a Yang-Mills connection. ∎
@@
Suppose $(A,B)$ is a cone-flat, i.e. a zero point of the cone Yang-Mills functional. If $\zeta$ is a harmonic form, then $A$ is a Yang-Mills connection.
@@
$d_{A}*F_{A}=-d_{A}[*\zeta B]=-(d*\zeta)B-(-1)^{m-2}(*\zeta)d_{A}B=0\,,$
@@
When $(M^{2n},\omega)$ is a symplectic manifold and $\zeta=\omega$, then a connection $A$ satisfying the curvature condition $F_{A}=-\omega B$ is called a symplectically-flat connection as introduced in [TZ]. Symplectically-flat connections are Yang-Mills connections with respect to a compatible metric. With such a metric, $*\omega=\omega^{n-1}/(n-1)!$, and therefore, $\omega$ is a harmonic form. We note that $(A,B)$ is cone-flat with respect to $\zeta=\omega$ is equivalent to $A$ being symplectically-flat.
@@
Thus far, we have described special cone Yang-Mills solutions pair $(A,B)$ where the connection part $A$ is Yang-Mills. Such is not the generic case. Below, we shall give a simple cone-flat solution $(A,B)$ where $A$ is not Yang-Mills. In order not to contradict Lemma 2.3, $\zeta$, as in the example below, though $d$-closed, will not be harmonic.
@@
Let $E=Ad\,P$ be the associated adjoint bundle over a Riemann surface $\Sigma$ and $A^{\prime}$ be a non-flat Yang-Mills connection on $\Sigma$. Let $\zeta$ be the volume form of $\Sigma$ normalized such that the total volume of $\Sigma$ is one. Define $M$ to be the three-dimensional circle bundle $\pi:M\to\Sigma$ with Euler class given by $\zeta$. Since $A^{\prime}$ is Yang-Mills, its curvature which we can write as $F_{A^{\prime}}=\zeta\Phi$ satisfies $d_{A^{\prime}}\Phi=0$. For simplicity, we will also use $\zeta$, $A^{\prime}$ and $\Phi$ to denote their pullbacks on $M$.
Let $c\in\mathbb{R}$ and $\theta\in\Omega^{1}(M)$ be the global angular one-form of the circle bundle $M$, i.e. $d\theta=\zeta$. For $A=A^{\prime}+c\,\theta\Phi$ and $B=-(1+c)\Phi$ on the pullback bundle $\pi^{*}E$, we find
@@
The curvature on $M$ is then given by
@@
Thus, $F_{A}$ satisfy the abelian Yang-Mills equation. However, there does not exist a function $B$ on $T^{4}$ such that $(A,B)$ satisfy the cone Yang-Mills equations (2.9)-(2.10). Equation (2.9) implies in the abelian case $d^{*}(F_{A}+\zeta B)=0\,$. But since we know already $d^{*}F_{A}=0\,$, this means $B$ is required to satisfy

### v2

@@ 1 Introduction
For the physical motivation, the cone Yang-Mills functional can be considered as the dimensional reduction of the standard Yang-Mills functional over the $S^{1}$ fiber of a circle bundle, $\pi:X\to M$. Unlike a product space, a circle bundle generically is one where the fiber circle is non-trivially twisted over $M$. Such arises for instance in Kaluza-Klein monopole solutions with a non-vanishing background two-form flux $\zeta$. On any circle bundle $X$, there exists what is called a global angular one-form $\theta$ which can be locally expressed as $\theta=dz+a$ where $z$ denotes the $S^{1}$ fiber coordinate and $a$ is the $U(1)$ Kaluza-Klein gauge field. The two-form flux, $\zeta=d\theta=d(dz+a)=da$, is then the background field strength of the $U(1)$ gauge field and effectively measures the twisting of the $S^{1}$ fiber.
Now for the dimensional reduction of the Yang-Mills functional over the $S^{1}$ fiber of the circle bundle $\pi:X\to M$, we take the metric to be
@@
and further require that the local connection one-form (Yang-Mills gauge field) $\mathcal{A}$ on $X$ be invariant under translation of the circle fiber. Such a connection form on $X$ can be expressed as
@@
Here, both $A$ and $B$ takes values on $Ad\,P$, and $B$ represents the component of the connection in the circle direction. Since $\mathcal{A}$ is invariant under $S^{1}$ translation, both $A$ and $B$ only have dependence on the coordinates of $M$. The curvature two-form of $\mathcal{A}$ (Yang-Mills field strength) on $X$ then has the following expression:
@@
where we have used $d\theta=\zeta$. Comparing with the cone Yang-Mills functional in (1), we see that $S_{cYM}=\|\mathcal{F}_{\mathcal{A}}\|^{2}\,$, and therefore, it is just the Yang-Mills functional on $X$ dimensionally reduced to $M$ over the fiber circle. And we can also interpret the cone Yang-Mills Euler-Lagrange equations (1.2)-(1.3) as the dimensionally-reduced Yang-Mills equations over the $S^{1}$ fiber of $X$.
@@
where we have used a simplified notation, e.g. $F_{A}\wedge*F_{A}=|F_{A}|^{2}dV$. The action thus consists of a Yang-Mills part, an adjoint-valued scalar with a “mass” $|\zeta|$ (which may vary over $M$), and an interaction term between $A$ and $B$. Heuristically, the Euler-Lagrange equations of motion (1.2)-(1.3) can be thought of as a mixing of the Yang-Mills equations with a Klein-Gordon-type scalar plus an interaction term.
The cone Yang-Mills functional also has a mathematical motivation. In fact, the functional arises naturally from its relation with a mapping cone complex which we will provide an explanation here. (For a reference on the general mapping cone complex, see for example [Weibel] and also [CTT].)
@@
Since the grading of the two components of $\operatorname{Cone}^{k}(\zeta)$ in (1.8) are different, it is useful to introduce a formal one-form $\theta$ to supplement the second component so that we can express a cone form as a sum with both components having the same total degree $k$, i.e.
@@
Additionally, it is useful for us to impose that the formal one-form $\theta$ satisfy $d\theta=\zeta$. For this will allow us to interpret the cone differential $d_{C}$ simply as the exterior derivative:
@@
where $\mathcal{A}=A+\theta\wedge B\,$, with $A\in\Omega^{1}(M,Ad\,P)$ and $B\in\Omega^{0}(M,Ad\,P)$. The cone curvature then takes the form
@@
The cone Yang-Mills functional of (1) is then just the normed square of the cone curvature, $\|\mathcal{F}_{\mathcal{A}}\|^{2}$. It is worthwhile to emphasize that this mapping cone perspective only requires that the two-form $\zeta\in\Omega^{2}(M)$ be $d$-closed and nothing more. This is in contrast with the dimensional reduction perspective, where $\zeta$ mathematically represents the Euler class of the circle bundle $\pi:X\to M$ and hence would need to be an element of $H^{2}(M,\mathbb{Z})$.
In this paper, we take an important first step in understanding cone Yang-Mills solutions. We will show that a subset of the solutions involves Yang-Mills connections. For instance, in dimension two when $M$ is compact and $\zeta$ is taken to be the volume form, the above cone-flat condition (1.11) implies exactly the two-dimensional Yang-Mills condition
@@
On the other hand, it should be evident that the space of cone Yang-Mills solutions $(A,B)$ is generally much richer and different from that of Yang-Mills solutions. As we will show in explicit examples, not all Yang-Mills connection $A$ can be paired with a scalar $B$ to form a cone Yang-Mills solutions. Conversely, there are also cone Yang-Mills solutions $(A,B)$ where $A$ is not Yang-Mills. Furthermore, given a cone Yang-Mills solutions $(A,B)$, there may be other scalars $B^{\prime}$ such that $(A,B^{\prime})$ remain cone Yang-Mills.
Interestingly, when $M$ is three-dimensional, we are able to write down a Bogomolny-type condition, that gives a sufficient condition on $(A,B)$ to be a cone Yang-Mills solution. Such a condition can be motivated by considering $\mathcal{F}_{\mathcal{A}}$ as the Yang-Mills curvature on a four-dimensional circle bundle $X$ and imposing the (anti-)self-dual condition. Then, dimensionally reducing on the fiber $S^{1}$ by expressing $\mathcal{F}_{\mathcal{A}}$ as in (1.10), the (anti-)self-dual condition implies the three-dimensional condition
$F_{A}+\zeta B=\pm*d_{A}B\,.$ (1.13)
Notice that if we set $\zeta=0$, the above equation is just the Bogomolny monopole equation. Similar to the (anti-)self-dual Yang-Mills condition in four dimensions, (1.13) is a first-order condition whose solutions solve (1.2)-(1.3) in three dimensions. This duality condition will assist us in finding non-abelian, cone Yang-Mills solutions. Indeed, we shall give in Section 3.2 an explicit $SU(2)$ solution of (1.13) that comes from dimensionally reducing the four-dimensional Taub-NUT gravitational instanton solution.
@@
This paper is organized as follows. In Section 2, after a brief description of our notations/conventions, we proceed to consider the first-order variation of the cone Yang-Mills functional to obtain its Euler-Lagrange equations. We also show that modulo gauge equivalence, the cone Yang-Mills equations are elliptic and hence has a finite-dimensional solution space on a closed manifold. We also describe properties of cone Yang-Mills solutions under certain conditions for the two-form $\zeta$ and structure group $G$, and especially emphasizing its relationship to Yang-Mills connections. In Section 3, we work out the special case of abelian cone Yang-Mills solutions in dimension two. We also discuss the three-dimensional Bogomolny-type monopole condition (1.13) and give an explicit non-trivial $SU(2)$ cone Yang-Mills solutions on $M=\mathbb{R}^{3}-\{0\}$, i.e. the Euclidean space with the origin removed. Finally, in Section 4, we consider bundles that can carry cone-flat solutions when $\zeta$ is a non-degenerate, closed two-form, and prove the classification of cone-flat bundles of Theorem 1.1.
@@ 2.1 Preliminaries
Let $M$ be a smooth manifold and $P$ be a principal bundle over $M$. We consider the associated adjoint bundle $Ad\,P$ equipped with an inner product that is invariant under the adjoint action. Hence, for arbitrary $x,y,z\in\Omega^{*}(M,Ad\,P)$,
@@
Having defined the inner product on $\operatorname{Cone}^{*}(\zeta)(M,Ad\,P)$, we write down the Yang-Mills functional associated with the cone connection. Given a connection $A\in\Omega^{1}(M,Ad\,P)$ and a section $B\in\Omega^{0}(M,Ad\,P)$, we write the cone connection as $\mathcal{A}=A+\theta B\in\operatorname{Cone}^{1}(\zeta)(M,Ad\,P)$. Locally, if we write the usual covariant derivative as $d_{A}=d+A$, then the cone covariant derivative is given by $D_{\mathcal{C}}=d+\mathcal{A}=d+(A+\theta B)$ and the cone curvature takes the form
$\mathcal{F}_{\mathcal{A}}=D_{\mathcal{C}}D_{\mathcal{C}}$ $=d(A+\theta B)+(A+\theta B)\wedge(A+\theta B)$
@@
$S_{cY\!M}(A+\theta B)$ $=\|\mathcal{F}_{\mathcal{A}}\|_{\mathcal{C}}^{2}=\|(F_{A}+\zeta B)-\theta\,d_{A}B\|^{2}_{C}$
@@
$=\|\mathcal{F}_{\mathcal{A}}\|_{\mathcal{C}}^{2}+2t\big{(}\langle d_{A}^{*}(F_{A}+\zeta B)+[B,d_{A}B],\eta\rangle+\langle\zeta^{*}(F_{A}+\zeta B)+d_{A}^{*}d_{A}B,\xi\rangle\big{)}+o(t)$
@@ 2.2 Elliptic property of the cone Yang-Mills solutions
We can write (2.13)-(2.14) more simply in terms of $D_{\mathcal{C}}$, the cone covariant derivative. Locally, if $d_{A}=d+A$, then $D_{\mathcal{C}}=d+(A+\theta B)$ and
@@
The last equation follows from the relation $(-1)^{m+1}*[\xi,*d_{A}B]=[\xi,d_{A}B]$. This can be seen by writing $d_{A}B=\sum\mu_{i}\otimes\alpha_{i}$ and $\xi=\sum\xi_{i}\otimes\alpha_{i}$ where $\mu_{i}\in\Omega^{1}(M)$, $\xi_{i}\in\Omega^{0}(M)$ and $\{\alpha_{i}\}$ is a basis of the Lie algebra $\mathfrak{g}=\operatorname{Lie}(G)$. Then
@@
2.3 Comparison with Yang-Mills solutions
@@
Hence, cone Yang-Mills solutions with $B=0$ are a subset of Yang-Mills solutions satisfying additionally the second condition of (2.20).
In particular, when $M$ is a Kähler manifold and $\zeta=\omega$ is the Kähler metric, a hermitian Yang-Mills connection that satisfies the conditions
@@
is also a cone Yang-Mills solution satisfying (2.20) with $B=0$. This is because $\omega^{*}F_{A}=0$ is equivalent to $\omega^{n-1}\wedge F_{A}=0$ for a Kähler metric.
@@
When $\zeta=0$, $\zeta^{*}$ is a zero map. The second equation of cone Yang-Mills condition (2.10) becomes just $d_{A}^{*}d_{A}B=0$, which implies $d_{A}B=0$ on a compact manifold. The first equation (2.9) then simplifies to $d_{A}^{*}F_{A}=0$. Hence, $A$ must be a Yang-Mills connection with $B$ covariantly constant. ∎
@@
Suppose $(A,B)$ is a cone-flat solution, i.e. a zero point of the cone Yang-Mills functional. If $\zeta$ is a harmonic form, then $A$ is a Yang-Mills connection.
@@
$d_{A}*F_{A}=-d_{A}(*\,\zeta B)=-(d*\zeta)B-(-1)^{m-2}(*\zeta)d_{A}B=0\,,$
@@
When $(M^{2n},\omega)$ is a symplectic manifold and $\zeta=\omega$, then a connection satisfying the curvature condition $F_{A}=-\omega B$ such that $d_{A}B=0$ is called a symplectically-flat connection as introduced in [TZ]. Symplectically-flat connections are Yang-Mills connections with respect to a compatible metric. This agrees with Lemma 2.3 above since with respect to a compatible metric, $*\,\omega=\omega^{n-1}/(n-1)!\,$ which implies $d^{*}\omega=0\,$, i.e. $\omega$ is a harmonic form.
@@
Thus far, we have described special cone Yang-Mills solution pairs $(A,B)$ where the connection part $A$ is Yang-Mills. Such is not the generic case. Below, we shall give a simple cone-flat solution $(A,B)$ where $A$ is not Yang-Mills. In order not to contradict Lemma 2.3, the $\zeta$ in the example below is not harmonic.
@@
Let $\Sigma$ be a Riemann surface and $A^{\prime}\in\Omega^{1}(\Sigma,Ad\,P)$ be a non-flat Yang-Mills connection on $\Sigma$. Let $\zeta$ be the volume form of $\Sigma$ normalized such that the total volume of $\Sigma$ is one. Define $M$ to be the three-dimensional circle bundle $\pi:M\to\Sigma$ with Euler class given by $\zeta$. Since $A^{\prime}$ is Yang-Mills, we can write its curvature as $F_{A^{\prime}}=\zeta\Phi$ such that $d_{A^{\prime}}\Phi=0$. For simplicity, we will also use $\zeta$, $A^{\prime}$ and $\Phi$ to denote their pullbacks on $M$.
Let $\theta\in\Omega^{1}(M)$ be the global angular one-form of the circle bundle $M$, i.e. $d\theta=\zeta$, and also let $c\in\mathbb{R}$. For $A=A^{\prime}+c\,\theta\Phi$ and $B=-(1+c)\Phi$ on the pullback bundle $\pi^{*}(Ad\,P)$, we find
@@
The curvature two-form is then given by
@@
Thus, $F_{A}$ satisfies the abelian Yang-Mills equation. However, there does not exist a function $B$ on $T^{4}$ such that $(A,B)$ satisfy the cone Yang-Mills equations (2.9)-(2.10). Equation (2.9) implies in the abelian case $d^{*}(F_{A}+\zeta B)=0\,$. But since we know already $d^{*}F_{A}=0\,$, this means $B$ is required to satisfy